# Relativistic unitary coupled cluster method for ground-state molecular properties

### v1

@@ III Computational and implementation details
The four-component UCC3 and qUCCSD expectation value approach for the calculation of the first-order property has been implemented in our in-house quantum chemistry software package, BAGH.[60]
It is primarily developed in Python, with computationally intensive parts written in Fortran and Cython. BAGH relies on external software packages to generate one and two-electron integrals, and it is interfaced with PySCF[61, 62, 63], socutils[64], DIRAC[65], and GAMESS-US[66]. The DIRAC package is used to solve the DHF equations and to compute one- and two-electron integrals, along with one-electron property integrals for the PDM and HFS calculations. For the EFG calculation, the DHF orbital and the necessary integrals are instead generated using the PySCF package[61, 62, 63] for this work.
Dipole moment calculations were performed for a series of alkaline earth metal monofluoride (AF, where A = Mg, Ca, Sr, Ba), employing a quadruple-zeta basis set, Dyall.cv4z[67, 68] for the metal atoms and cc-pCVQZ[69] for the fluorine atom.
For HFS constant calculations, the aug-cc-pCV5Z[69] basis set was used for the hydrogen atom, aug-cc-pCVQZ for F, Be, and Mg atoms, and Dyall.cv4z for the Ca atom. For the MgH and MgF, a virtual orbital energy cutoff of 10 a.u. was applied, while a cutoff of 15 a.u. was used for CaH and MgF.
EFG calculations were carried out using Dyall.cv2z[70] basis set for all FX [where, X= F, Cl, Br, I, At] molecules.
@@ IV.1 Permanent Dipole Moment
Table 1 summarizes the PDM of the Group II monofluoride molecules (MgF, CaF, SrF, and BaF) in UCC3, qUCCSD, and CCSD Z-vector methods, along with their available experimental data. All the calculated values are presented in Debye units. To ensure consistency with the Z-vector method, we computed our results using the same molecular geometry, basis set, and cut-off energy as those employed by Pal and co-workers.[71]
Figure 1(a) presents the comparison of UCC3, the qUCCSD method, with the CCSD Z-vector method, and it can be seen that the qUCCSD results show an excellent agreement with the CCSD Z-vector approach. Whereas the deviation in the UCC3 method is quite high. This discrepancy arises because the qUCCSD method contains a more complete treatment of the single excitation (up to quadratic term in the amplitudes equation), whereas they are linear in the UCC3 amplitude equation. For the MgF molecule, both UCC3 and qUCCSD approaches agree remarkably with each other, but the agreement deteriorates as we go down the group. In the case of CaF, the UCC3 method shows a deviation of 0.22 Debye from the CCSD Z-vector results, while the deviation from the CCSD Z-vector results is 0.01 Debye for the qUCCSD method. Similarly, for SrF, the deviations from the CCSD Z-vector method are 0.39 Debye and 0.03 Debye for UCC3 and qUCCSD, respectively. The BaF molecule shows the largest deviation for UCC3 at 0.82 Debye, whereas the qUCCSD method maintains high accuracy with a deviation of just 0.01 Debye. These findings demonstrate that the qUCCSD method yields results in excellent agreement with those obtained using the Z-vector approach.
We have compared our results with experimental measurements for CaF, SrF, and BaF. No experimental measurements are available for MgF. To assess the accuracy of our results, we compared them with available experimental data, with deviations expressed as percentage errors, denoted by .
Figure 1(b) presents a comparison of from the experimental results for UCC3, qUCCSD, and CCSD Z-vector methods. It can be seen that the UCC3 method displays a larger deviation from the experimental values, whereas the qUCCSD and Z-vector methods yield more accurate results and are generally very close to each other. The UCC3 method shows the smallest deviation with respect to the experiment for CaF and is 8.15 . The qUCCSD and Z-vector methods yield much lower deviations of 1.70 and 1.30, respectively, and it is observed for SrF. The UCC3 shows the highest deviation of 22.63 for BaF. The qUCCSD and Z-vector results show very good agreement with the experiment for BaF with remarkably improved values of 3.36, and 3.11, respectively.
@@ IV.2 Hyperfine Structure Constant
The HFS constants are sensitive to the quality of the wave function in the nuclear region and can be a good test for the quality of the wave function obtained from a particular method. For this purpose, we have computed the HFS constants for a series of diatomic molecules, including BeH, MgH, CaH, and CaF, using both the UCC3 and qUCCSD methods. The computed parallel () and perpendicular () components of the HFS constants are presented in Tables 2 and 3 , respectively. To ensure consistency, our results are compared with those from the Z-vector method, using identical basis sets, geometries, and cutoff values as described in the reference.[75] The data presented in Tables 2 and 3 reveal that the qUCCSD results show better agreement with both Z-vector and the experimental values compared to the UCC3 results, which exhibit more deviation with respect to the experiment. The maximum and minimum deviation of UCC3 and qUCCSD for values from the Z-vector results are observed in different systems. For UCC3, the largest deviation (92.87 MHz) from the CCSD Z vector result occurs in 1H in MgH, while the smallest (0.82 MHz) is found for 25Mg in MgF. In contrast, qUCCSD shows a maximum deviation of 6.61 MHz for 19F in MgF and a minimum of 0.89 MHz for 1H in CaH.
@@
In the case of , the maximum deviation of 93.64 MHz from the CCSD Z-vector method is obtained in UCC3 again for 1H in MgH. The minimum deviation for UCC3 (0.63 MHz) is observed for 25Mg in MgF. The maximum deviation in the qUCCSD method from the CCSD Z-vector is 4.98 MHz observed for 25Mg in MgF. Figures 3 (a) and 3 (b) present a comparison of the and values, respectively, obtained using UCC3, qUCCSD and Z-vector method. From the above observation, the results clearly demonstrate that the qUCCSD approach consistently yields better agreement with the CCSD Z-vector results compared to the UCC3 method. Our calculated results have also been compared with the available experimental data. In figures 3(a) and 3(b) represent the percentage deviation () of the and values, respectively, obtained using UCC3, qUCCSD, and Z-vector method. For , the largest in UCC3 occurs at 1H in CaH, which 51.34 relative deviation from experimental. The higher relative deviation observed in the qUCCSD and CCSD Z-vector methods is 17.42 MHz and 16.77 MHz, respectively, and it is observed for the 25 Mg nucleus in MgH. In the case of the component, the highest deviation is observed for the 1H nucleus in the CaH molecule, and it shows a 51.22 with the UCC3 method. The qUCCSD approach shows its highest deviation of 23.78 for the 19F nucleus in CaF, which is close to the corresponding CCSD Z-vector value of 20.94 . From the overall trends observed in the table, it reveals that the qUCCSD method demonstrates very good agreement with experimental values for both and of the hyperfine coupling constant, showing close agreement with the corresponding Z-vector results. The UCC3 method shows significant deviation with respect to experiments in some cases.
@@ IV.3 Electric Field Gradient
To further assess the performance of the newly developed UCC methods, we have calculated the electronic component of the EFG value () for the series of dihalogen molecules FY (Y = F, Cl, Br, I, At) and compared them with the four-component CCSD Z-vector results. The reference CCSD Z-vector values were adopted from the work of Aucar et al.[79] To maintain consistency, all calculations in this work were carried out using the same dyall.cv2z basis set and molecular geometries as employed in the ref [79]. The molecular geometries for the series of dihalogen molecules are also provided in the supporting information. Figure 4(a) provides a comparison of the UCC3 and qUCCSD values with the reference CCSD Z-vector method for the zz-component (with the z axis chosen along the molecular bond) of the electronic component of EFG values () for atoms in the FY series. The corresponding values are provided in Table 4. It can be seen that the qUCCSD values closely follow the CCSD values throughout the data set, with the bars almost overlapping in most cases. The UCC3 method displays small but noticeable deviations, particularly for heavier atoms such as I and At. Table 4 also presents their deviations from the reference values, which in this case are CCSD Z-vector values, as no experimental values are available. From the table, it is clear that qUCCSD values show smaller deviations from CCSD results compared to UCC3, often by an order of magnitude. The highest deviation in UCC3 is observed to be 1.008 a.u., which is observed for At in FAt. The qUCCSD method, on the other hand, shows a deviation of 0.086 a.u. for the electronic component of EFG values ().
@@ V Conclusion
We present the theory, implementation, and benchmarking of the first-order property calculation using the unitary coupled cluster expectation value method. The Hermitian formulation within the qUCCSD method allows a simple definition of the first-order property using the expectation value approach and does not require the calculation of an additional set of amplitudes, as in the standard coupled cluster Z-vector method. We have calculated the permanent dipole moment, hyperfine structure constant, and electric field gradient of a series of molecules containing heavy atoms. The calculated values using the qUCCSD expectation value approach are in close agreement with the CCSD Z-vector results and are consistent with the available experimental values. However, the perturbative UCC3 method does not give consistent results and often shows large errors. The cause is presumably the incomplete treatment of the single excitation in the UCC3 level. The present work shows that the UCC method gives an attractive option for the calculation of first-order properties even on a classical computer. It would be interesting to extend the unitary coupled cluster method to second and higher-order properties. Work is in progress in that direction.
@@
The Supplementary Material contains the programmable expressions for the ground state first-order property in the UCC3 and qUCCSD methods and the molecular geometry for the series of dihalogen molecules.
@@ Acknowledgements.
The authors acknowledge the support from the IIT Bombay, DST-SERB CRG (Project No. CRG/2022/005672) and MATRICS (Project No. MTR/2021/000420) projects, and IIT Bombay supercomputational facility and C-DAC supercomputing resources (PARAM Smriti and PARAM Brahma) for computational time. KM acknowledges CSIR-HRDG for a senior research fellowship. SC acknowledges the Prime Minister’s Research Fellowship (PMRF). The authors acknowledge Juan J. Aucar (Institute for Modelling and Innovative Technology, IMIT (CONICET-UNNE), Argentina) for his valuable feedback and discussions during the preparation of this manuscript.References
@@
- Visscher, Lee, and Dyall [1996a] L. Visscher, T. J. Lee, and K. G. Dyall, J. Chem. Phys. 105, 8769 (1996a).
@@
- Reiher and Wolf [2014] M. Reiher and A. Wolf, Relativistic quantum chemistry: the fundamental theory of molecular science (John Wiley & Sons, 2014).
@@
- Shavitt and Bartlett [2009] I. Shavitt and R. J. Bartlett, Many-body methods in chemistry and physics: MBPT and coupled-cluster theory (Cambridge university press, 2009).
@@
- Visscher, Lee, and Dyall [1996b] L. Visscher, T. J. Lee, and K. G. Dyall, J. Chem. Phys. 105, 8769 (1996b).
- Petrov et al. [2002] A. Petrov, N. Mosyagin, T. Isaev, A. Titov, V. Ezhov, E. Eliav, and U. Kaldor, Phys. Rev. Lett. 88, 073001 (2002).
@@
- Prasannaa et al. [2016] V. Prasannaa, S. Sreerekha, M. Abe, V. Bannur, and B. Das, Phys. Rev. A 93, 042504 (2016).
@@
- Das and Mukherjee [2005] B. Das and D. Mukherjee, Eur. Phys. J. D 32, 25 (2005).
@@
- Hodecker et al. [2020] M. Hodecker, S. M. Thielen, J. Liu, D. R. Rehn, and A. Dreuw, J. Chem. Theory Comput. 16, 3654 (2020), pMID: 32396348, https://doi.org/10.1021/acs.jctc.0c00335 .
@@
- Majee et al. [2025] K. Majee, J. Šimunek, J. Noga, and A. K. Dutta, “A perturbative triples correction to relativistic quadratic unitary coupled cluster method: Theory, implementation and benchmarking,” (2025), arXiv:2509.03634 [physics.chem-ph] .
@@
- Dutta et al. [2023] A. Dutta, A. Manna, B. Jangid, K. Majee, K. Surjuse, M. Mukherjee, M. Thapa, S. Arora, S. Chamoli, S. Haldar, et al., “Bagh: A quantum chemistry software package,” (2023).
- Sun et al. [2020] Q. Sun, X. Zhang, S. Banerjee, P. Bao, M. Barbry, N. S. Blunt, N. A. Bogdanov, G. H. Booth, J. Chen, Z.-H. Cui, et al., J. Chem. Phys. 153 (2020).
- Sun et al. [2018] Q. Sun, T. C. Berkelbach, N. S. Blunt, G. H. Booth, S. Guo, Z. Li, J. Liu, J. D. McClain, E. R. Sayfutyarova, S. Sharma, et al., Wiley Interdiscip. Rev. Comput. Mol. Sci. 8, e1340 (2018).
@@
- Jensen et al. [2022] H. A. Jensen, R. Bast, A. Gomes, T. Saue, L. Visscher, I. Aucar, V. Bakken, C. Chibueze, J. Creutzberg, and K. Dyall, (2022).
- Barca et al. [2020] G. M. Barca, C. Bertoni, L. Carrington, D. Datta, N. De Silva, J. E. Deustua, D. G. Fedorov, J. R. Gour, A. O. Gunina, E. Guidez, et al., J. Chem. Phys. 152 (2020).
@@
- Anand et al. [2022] A. Anand, P. Schleich, S. Alperin-Lea, P. W. Jensen, S. Sim, M. Díaz-Tinoco, J. S. Kottmann, M. Degroote, A. F. Izmaylov, and A. Aspuru-Guzik, Chem. Soc. Rev. 51, 1659 (2022).

### v2

@@ III Computational and implementation details
The four-component UCC3 and qUCCSD expectation-value approaches for the calculation of first-order properties have been implemented in our in-house quantum chemistry software package, BAGH.[61]
The package is primarily developed in Python, with computationally intensive components written in Fortran and Cython. BAGH relies on external software packages for the generation of one- and two-electron integrals and is interfaced with PySCF[62, 63, 64], socutils,[65] DIRAC,[66] and GAMESS-US.[67] The DIRAC package is used to solve the DHF equations and to compute one- and two-electron integrals, along with one-electron property integrals required for the PDM and HFS calculations. For the EFG calculations, the DHF orbitals and the required integrals are generated using the PySCF package.[62, 63, 64]
Dipole moment calculations were performed for a series of alkaline earth metal monofluoride (AF, where A = Mg, Ca, Sr, Ba), employing a quadruple-zeta basis set: Dyall.cv4z[68, 69] for the metal atoms and cc-pCVQZ[70] for fluorine.
For HFS constant calculations, the aug-cc-pCV5Z[70] basis set was used for hydrogen, aug-cc-pCVQZ for F, Be, and Mg, and Dyall.cv4z for Ca.
EFG calculations were carried out using Dyall.cv2z[71] basis set for all FX molecules[where, X= F, Cl, Br, I, At].
@@ IV.1 Permanent Dipole Moment
Table 1 summarizes the permanent dipole moments (PDMs) of the Group II monofluoride molecules (MgF, CaF, SrF, and BaF) obtained using the UCC3, qUCCSD, and Z-vector methods, along with available experimental data. All calculated values are reported in Debye. To ensure consistency with the available Z-vector results, our calculations were performed using the same molecular geometries, basis sets, and cutoff energies as those employed by Pal and co-workers.[72]
Figure 1(a) compares the UCC3 and qUCCSD results with those obtained using the Z-vector method and shows that the qUCCSD results are in excellent agreement with the Z-vector approach. In contrast, the deviation obtained with the UCC3 method is significantly larger. This discrepancy arises because the qUCCSD method includes a more complete treatment of single excitations (up to quadratic terms in the amplitude equations), whereas these contributions are linear in the UCC3 amplitude equations. For MgF, the UCC3 and qUCCSD approaches agree remarkably well with each other; however, this agreement deteriorates as one moves down the group. In the case of CaF, the UCC3 method shows a deviation of 0.22 Debye from the Z-vector result, whereas the corresponding deviation for the qUCCSD method is only 0.01 Debye. Similarly, for SrF, the deviations from the CCSD Z-vector method are 0.39 Debye and 0.03 Debye for the UCC3 and qUCCSD, respectively. For BaF, the UCC3 method shows the largest deviation, amounting to 0.82 Debye, whereas the qUCCSD method maintains high accuracy with a deviation of only 0.01 Debye. These findings demonstrate that the qUCCSD method yields results in excellent agreement with those obtained using the Z-vector approach.
We have compared our results with available experimental measurements for CaF, SrF, and BaF. No experimental measurements are available for MgF. To assess the accuracy of our results, deviations from the experimental values were expressed as percentage errors, denoted by as
Figure 1(b) compares the percentage deviations () of the UCC3, qUCCSD, and Z-vector results from the experimental values. It can be seen that the UCC3 method displays larger deviations from the experimental values, whereas the qUCCSD and Z-vector methods yield more accurate results and are generally very close to each other. For CaF, the UCC3 method shows the smallest deviation from experiment, amounting to 8.15 . For SrF, the qUCCSD and Z-vector methods yield much lower deviations of 1.70 and 1.30, respectively. The UCC3 method shows the highest deviation for BaF, amounting to 22.63. For BaF, the qUCCSD and Z-vector results show very good agreement with experiment, with markedly improved deviations of 3.36, and 3.11, respectively.
@@ IV.2 Hyperfine Structure Constant
Hyperfine structure (HFS) constants are sensitive to the quality of the wave function in the nuclear region and therefore provide a stringent test of the wave function obtained using a given method. Accordingly, we computed the HFS constants for a series of diatomic molecules, including BeH, MgH, CaH, and CaF, using both the UCC3 and qUCCSD methods.The computed parallel () and perpendicular () components of the HFS constants are presented in Tables 2 and 3, respectively. To ensure consistency, our results are compared with those obtained using the Z-vector method, employing identical basis sets, molecular geometries, and cutoff values as described in Ref.[76]. The data presented in Tables 2 and 3 reveal that the qUCCSD results show better agreement with both Z-vector and the experimental values than the UCC3 results, which exhibit larger deviations from experiment. The maximum and minimum deviations of UCC3 and qUCCSD for values from the Z-vector results occur for different systems. For UCC3, the largest deviation (92.87 MHz) from the Z-vector result occurs in 1H in MgH, while the smallest (0.82 MHz) is found for 25Mg in MgF. In contrast, the qUCCSD method shows a maximum deviation of 6.61 MHz for 19F for MgF and a minimum deviation of 0.89 MHz for 1H in CaH.
@@
In the case of , the maximum deviation of 93.64 MHz from the Z-vector method is obtained in UCC3 again for 1H in MgH. The minimum deviation obtained with the UCC3 method (0.63 MHz) is observed for 25Mg in MgF. For the qUCCSD method, the maximum deviation from the Z-vector result is 4.98 MHz, observed for 25Mg in MgF. Figures 2(a) and 2(b) present comparisons of the and values, respectively, obtained using the UCC3, qUCCSD, and Z-vector methods. The results clearly demonstrate that the qUCCSD method consistently yields better agreement with the Z-vector results than the UCC3 method. Our calculated results were also compared with available experimental data. Figures 3(a) and 3(b) show the percentage deviation () of the and values, respectively, obtained using UCC3, qUCCSD, and Z-vector methods. For the component, the largest for the UCC3 method occurs for 1H in CaH, corresponding to a relative deviation of 51.34 from experiment. The higher relative deviations observed in the qUCCSD and Z-vector methods are 17.42 and 16.77 , respectively, both occurring for the 25 Mg nucleus in MgH. In the case of the component, the largest deviation is observed for the 1H nucleus in CaH, amounting to 51.22 for the UCC3 method. The qUCCSD approach shows its highest deviation of 23.78 , for the 19F nucleus in CaF, which is close to the corresponding Z-vector value of 20.94 . From the overall trends observed in the table, it is evident that the qUCCSD method demonstrates very good agreement with experimental values for both and components of the hyperfine coupling constant, showing close agreement with the corresponding Z-vector results. In contrast, the UCC3 method exhibits significant deviations from experimental results in several cases.
@@ IV.3 Electric Field Gradient
To further assess the performance of the newly developed UCC methods, we calculated the electronic component () of the electric field gradient (), for a series of dihalogen molecules, FY (Y = F, Cl, Br, I, At), and compared them with the four-component CCSD results. The reference Z-vector values were taken from the work of Aucar et al..[80] To maintain consistency, all calculations were carried out using the same Dyall.cv2z basis set and molecular geometries as employed in Ref. 80. The molecular geometries for the series of dihalogen molecules are also provided in the Supporting Information. Figure 4(a) compares the UCC3 and qUCCSD values with the reference Z-vector results for the zz-component (with the z-axis chosen along the molecular bond) of the EFG values (), for atoms in the FY series. The corresponding numerical values are listed in Table 4. It can be seen that the qUCCSD values closely follow the CCSD results throughout the data set, with the bars nearly overlapping in most cases. In contrast, the UCC3 method displays small but noticeable deviations, particularly for heavier atoms such as I and At. Table 4 also reports the deviations() from the reference values, which in this case correspond to the Z-vector results, as no experimental data are available. From Table 4, it is clear that the qUCCSD values show smaller deviations from the CCSD results than the UCC3 values, often by an order of magnitude. The largest deviation obtained with the UCC3 method is 1.008 a.u., observed for At in FAt. In contrast, the qUCCSD method shows a deviation of only 0.044 a.u. for the same system.
@@ V Conclusion
We present the theory, implementation, and benchmarking of first-order property calculations using the relativistic unitary coupled-cluster expectation-value method. The Hermitian formulation within the qUCCSD method allows a simple definition of first-order properties using the expectation-value approach and does not require the calculation of an additional set of amplitudes, as required in the standard coupled cluster Z-vector method. We calculated the permanent dipole moment, hyperfine structure constant, and electric field gradient for a series of molecules containing heavy atoms. The values obtained using the qUCCSD expectation-value approach are in close agreement with the CCSD Z-vector results and are consistent with available experimental data. However, the perturbative UCC3 method does not exhibit consistent performance and often shows large errors. This behavior can presumably be attributed to the incomplete treatment of single excitations at the UCC3 level. The present work demonstrates that the projection-based approximation to the UCC method provides an attractive option for the calculation of first-order properties, even on a classical computer. It would be interesting to extend the unitary coupled cluster method to second- and higher-order properties. Work in this direction is currently in progress.
@@
The Supplementary Material contains the programmable expressions for the ground state first-order property in the UCC3 and qUCCSD methods, the derivation of the equivalence of the energy derivative and expectation value approach in UCC at the full CI limit, the molecular geometries for the series of dihalogen molecules, and the nuclear spins and magnetic moments of isotopes used in the calculations of the hyperfine structure constant.
@@ Acknowledgements.
AKD, SC, and KM acknowledge the support from IIT Bombay, ANRF(CRG/2023/002558), and ISRO for financial support. The authors also acknowledge the IIT Bombay supercomputing facility and C-DAC (Param Smriti and Param Bramha) for providing computational resources. SC acknowledges the Prime Minister’s Research Fellowship (PMRF), and KM acknowledges CSIR-HRDG for the research fellowship. AKD acknowledges the research fellowship funded by the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under project No. 09I03-03-V04-00117. The authors acknowledge Juan J. Aucar (Institute for Modelling and Innovative Technology, IMIT (CONICET-UNNE), Argentina) for his valuable feedback and discussions on EFG calculations during the preparation of this manuscript.References
@@
- Visscher, Lee, and Dyall [1996] L. Visscher, T. J. Lee, and K. G. Dyall, J. Chem. Phys. 105, 8769 (1996).
@@
- Reiher and Wolf [2014] M. Reiher and A. Wolf, Relativistic Quantum Chemistry: The Fundamental Theory of Molecular Science, 2nd ed. (Wiley-VCH, Weinheim, 2014).
@@
- Shavitt and Bartlett [2009] I. Shavitt and R. J. Bartlett, Many-body methods in chemistry and physics: MBPT and coupled-cluster theory (Cambridge University Presss, 2009).
@@
- Petrov et al. [2002] A. N. Petrov, N. S. Mosyagin, T. A. Isaev, A. V. Titov, V. F. Ezhov, E. Eliav, and U. Kaldor, Phys. Rev. Lett. 88, 073001 (2002).
@@
- Prasannaa et al. [2016] V. S. Prasannaa, S. Sreerekha, M. Abe, V. M. Bannur, and B. P. Das, Phys. Rev. A 93, 042504 (2016).
@@
- Das and Mukherjee [2005] B. P. Das and D. Mukherjee, Eur. Phys. J. D 32, 25 (2005).
@@
- Hodecker et al. [2020] M. Hodecker, S. M. Thielen, J. Liu, D. R. Rehn, and A. Dreuw, J. Chem. Theory Comput. 16, 3654 (2020), pMID: 32396348.
@@
- Majee et al. [2026] K. Majee, J. Šimunek, J. Noga, and A. K. Dutta, J. Chem. Phys. 164, 044109 (2026).
@@
- Brown and Carrington [2003] J. M. Brown and A. Carrington, Rotational Spectroscopy of Diatomic Molecules (Cambridge University Press, 2003).
- Frosch and Foley [1952] R. Frosch and R. M. Foley, Phys. Rev. 88, 243 (1952).
- Dutta et al. [2025] A. K. Dutta, A. Manna, B. Jangid, K. Majee, K. Surjuse, M. Mukherjee, M. Thapa, S. Arora, S. Chamoli, S. Haldar, S. Chakraborty, S. Mandal, and T. Mukhopadhyay, “BAGH: A quantum chemistry software package,” (2025).
- Sun et al. [2020] Q. Sun, X. Zhang, S. Banerjee, P. Bao, M. Barbry, N. S. Blunt, N. A. Bogdanov, G. H. Booth, J. Chen, Z.-H. Cui, J. J. Eriksen, Y. Gao, S. Guo, J. Hermann, M. R. Hermes, K. Koh, P. Koval, S. Lehtola, Z. Li, J. Liu, N. Mardirossian, J. D. McClain, M. Motta, B. Mussard, H. Q. Pham, A. Pulkin, W. Purwanto, P. J. Robinson, E. Ronca, E. R. Sayfutyarova, M. Scheurer, H. F. Schurkus, J. E. T. Smith, C. Sun, S.-N. Sun, S. Upadhyay, L. K. Wagner, X. Wang, A. White, J. D. Whitfield, M. J. Williamson, S. Wouters, J. Yang, J. M. Yu, T. Zhu, T. C. Berkelbach, S. Sharma, A. Y. Sokolov, and G. K.-L. Chan, “Recent developments in the PySCF program package,” (2020).
- Sun et al. [2018] Q. Sun, T. C. Berkelbach, N. S. Blunt, G. H. Booth, S. Guo, Z. Li, J. Liu, J. D. McClain, E. R. Sayfutyarova, S. Sharma, S. Wouters, and G. K.-L. Chan, “PySCF: The Python-based simulations of chemistry framework,” (2018).
@@
- Saue et al. [2025] T. Saue, L. Visscher, H. J. A. Jensen, R. Bast, A. Severo Pereira Gomes, I. A. Aucar, V. Bakken, J. Brandejs, C. Chibueze, J. Creutzberg, K. G. Dyall, S. Dubillard, U. Ekström, E. Eliav, T. Enevoldsen, E. Fasshauer, T. Fleig, O. Fossgaard, K. J. Gaul, L. Halbert, E. D. Hedegård, T. Helgaker, B. Helmich-Paris, J. Henriksson, M. van Horn, M. Ilias, C. Jacob, S. Knecht, S. Komorovsky, O. Kullie, J. K. Laerdahl, C. V. Larsen, Y. S. Lee, N. H. List, H. S. Nataraj, M. K. Nayak, P. Norman, A. Nyvang, M. Olejniczak, J. Olsen, J. M. H. Olsen, A. Papadopoulos, Y. C. Park, J. K. Pedersen, M. Pernpointner, J. Pototschnig, R. Di Remigio Eikås, M. Repisky, C. M. R. Rocha, K. Ruud, P. Salek, B. Schimmelpfennig, B. SENJEAN, A. Shee, J. Sikkema, A. Sunaga, A. J. Thorvaldsen, J. Thyssen, J. van Stralen, M. López Vidal, s. villaume, O. Visser, T. Winther, S. Yamamoto, and X. Yuan, “Dirac25,” (2025).
- Barca et al. [2020] G. M. J. Barca, C. Bertoni, L. Carrington, D. Datta, N. De Silva, J. E. Deustua, D. G. Fedorov, J. R. Gour, A. O. Gunina, E. Guidez, T. Harville, S. Irle, J. Ivanic, K. Kowalski, S. S. Leang, H. Li, W. Li, J. J. Lutz, I. Magoulas, J. Mato, V. Mironov, H. Nakata, B. Q. Pham, P. Piecuch, D. Poole, S. R. Pruitt, A. P. Rendell, L. B. Roskop, K. Ruedenberg, T. Sattasathuchana, M. W. Schmidt, J. Shen, L. Slipchenko, M. Sosonkina, V. Sundriyal, A. Tiwari, J. L. Galvez Vallejo, B. Westheimer, M. Włoch, P. Xu, F. Zahariev, and M. S. Gordon, “Recent developments in the general atomic and molecular electronic structure system,” (2020).
@@
- Childs et al. [1984] W. J. Childs, L. S. Goodman, U. Nielsen, and V. Pfeufer, J. Chem. Phys. 80, 2283 (1984).